Given the following 2 constraints, which solution is a feasible solution for a maximization problem?
(1)       14x1 + 6x2 ≤ 42
(2)       x1x2 ≤ 3

Group of answer choices

a. (x1, x2 ) = (2,1)

b. (x1, x2 ) = (1,5)

c. (x1, x2 ) = (5,1)

d. (x1, x2 ) = (4,4)

e. (x1, x2 ) = (2,6)
Answer #1

We apply given values of X1,X2 to first formula:

The results as follows :

a)28+ 6 <= 42

b) 14 + 30 <=44

c) 70 + 6 <= 42

d) 56 + 24 <= 42

e) 28 + 36 <= 42

Out of above except a, others are not feasible.

Answer : option a